# Analyticity of the Black Hole S-Matrix

### v1

@@ 3.2 Inspection of the exact solution
As required by causality, the Green’s function is analytic in the UHP of $\omega$ for finite $x$ and $x^{\prime}$. However, as anticipated by our general proof of analyticity, $R(\omega)$ has poles on the positive imaginary axis where $A_{\mathrm{out}}(\omega)$ diverges, due to the term $\Gamma\left(\frac{i\omega}{\alpha}\right)$. These poles are spurious, as they lack physical interpretation in this toy model. Finally, we confirm that both $G_{R}(\omega,x,x^{\prime})$ and $R(\omega)$ are analytic within the strip $\alpha<\text{Im}\,\omega<\alpha$ due to the exponential decay of the potential (77) (as shown in Appendix A).
@@
We first notice that $\phi_{L}(\omega,x)$ in (78) is manifestly analytic at those points, because $_{2}F_{1}(a,b,c,\xi)$ only has singularities when $c$ is a non-positive integer. To see how the singularities of $R(\omega)$ appear, we rewrite the hypergeometric function using the identity
@@ 4.1 Analyticity domains of the S-matrix
Importantly, the spherical symmetry of the Schwarzschild background reduces the problem to one-dimensional scattering of waves in the radial direction $r$. However, since the gravitational potential behaves like $1/r$ at large distances, the asymptotic solutions are not given in terms of plane waves, and the S-matrix is not rigorously defined. Instead, the radial solutions reduce to Coulomb-type wavefunctions, which behave like
@@
the potential decays fast enough and the asymptotic states reduce to plane waves, allowing us to define an S-matrix as in (4). [^12] An analogous IR-safe definition exists for electromagnetic (Coulomb) scattering in relativistic quantum mechanics Lippstreu:2025jit.
On the opposite end, the tortoise coordinate maps the event horizon $r=R_{s}$ to $x=$-$\infty$, where waves can escape, leaving the system. Wave scattering on the Schwarzschild background therefore amounts to scattering on the real line with open boundary conditions at both ends, which will allow us to rigorously apply the results of Section 2.
@@
Figure 9: The Schwarzschild potential barrier $V_{\text{RW}}(r(x))$ in units $R_{s}=1$ with $\ell=1,s=0$ for Left: $x\in\mathbb{R}$ and Right: $x\in\mathbb{C}$. The potential has an infinite number of poles on the imaginary axis, located at $x=i(1+2n)\pi R_{s},\;\;n\in\mathbb{Z}$. Each pole is a branch point, with the branch cuts chosen to run to $i\infty$ ($i\infty$) for branch points in the UHP (LHP). The potential decays exponentially as $x\rightarrow$-$\infty$ and decays as $x^{-2}$ as $x\to\infty$. The solid line on the right plot is the real axis, and the dashed line is slated by an angle $\theta$. We can take $\theta\to\pi/2-\epsilon$ without encountering singularities.
@@
The Regge–Wheeler potential in the complex $x$-plane is shown in the right panel of Figure 9, confirming the analytic structure discussed above. [^14] Note that the region of the imaginary axis between $i\pi\,R_{s}$ and $i\pi\,R_{s}$ is free of branch cuts, so the potential remains analytic along this strip. The Zerilli potential exhibits the same analytic features and is therefore not shown here.
@@
Since the potentials are strictly positive, all poles lie in the lower half of the complex frequency plane (see Section 2.6). These poles correspond to quasinormal modes and contribute to the retarded Green’s function, as well as to the transmission and reflection amplitudes.
@@ 4.2 Low-frequency expansion
where the real part of the phase-shift $\delta_{\ell}(\omega)$ captures elastic scattering and the imaginary part encodes absorptive effects. The phase shifts for $\ell=0,1$, obtained from the MST method Mano:1996vt; Mano:1996gn; Mano:1996mf, and corrected for consistency with the reflection symmetry and analyticity (7) take the form
@@ 4.3 Inspection of solvable cases
Newtonian potential. In the large-$r$ limit, the Regge–Wheeler potential in (95) can be approximated as
@@
The solutions of Eq. (102) are the well-known Coulomb wavefunctions. Rather than constructing the retarded Green’s function from the Jost solutions $\phi_{L}(\omega,x)$ and $\phi_{R}(\omega,x)$, we instead employ $\widetilde{\phi}_{L}(\omega,r)$, defined as the solution that is regular and vanishes at the origin and $\widetilde{\phi}_{R}(\omega,r)\equiv\phi_{R}(\omega,r)$, which still represents an outgoing plane wave in the limit $x\to\infty$. Applying these boundary conditions, the solutions are given by
@@
and the reflection amplitude is given by, via (2),
@@
The retarded Green’s function is analytic in the UHP of $\omega$, since the confluent hypergeometric functions are analytic for the relevant parameter ranges. This is shown explicitly in the left panel of Figure 10. Furthermore, $\widetilde{\phi}_{L}(\omega,r)$ is meromorphic everywhere (with poles in LHP), while $\widetilde{\phi}_{R}(\omega,r)$, and in turn $G_{R}(\omega,r,r^{\prime})$, has a branch cut along the negative imaginary axis of $\omega$. The discontinuity across this branch cut contributes to the late-time power-law decay tail of the black hole ringdown Leaver:1986gd; price_nonspherical_1972_2439; price_nonspherical_1972_2419; Casals:2012ng; Casals:2015nja. [^18] This can be seen by computing the inverse Fourier transform of the retarded Green’s function and deforming the integration contour into the LHP. Picking up the discontinuity of the branch cut results in the polynomial suppression of $G_{R}(t\to\infty,r,r^{\prime})$.
Figure 10: Solution to Newtonian potential in units $R_{s}=1$ for $\ell=1$. Dotted lines mark the real and imaginary axes. Left: $G_{R}(\omega,r=2,r^{\prime}=2.5)$. Right: $R(\omega)$. The Green’s function is analytic in the UHP, with QNMs (poles) appearing in the LHP. In the small $\omega$ approximation, the QNMs lie on the imaginary axis. The reflection amplitude has QNMs in the LHP as well, but also has a branch cut along the entire imaginary axis arising from the Stokes phenomenon.
@@
The presence and logarithmic nature of the singularitiy are consistent with the low-frequency approximation of the previous section, where such logarithms are also present. Additionally, the location of the branch cut was anticipated from Section 2, since the structure of the potential implies that any branch cuts can only lie on the imaginary axis. We will see this same structure in the high-frequency approximation.
High-frequency approximation. We now approximate the Regge-Wheeler equation (92) in the high-frequency limit $|R_{s}\omega|\gg 1$. Following the derivation in liu_spectrum_1996; Andersson:1996cm, the Regge-Wheeler equation can be written as
@@
whose solutions are known. The centrifugal term in the potential is subleading in this limit, so at this order the solutions are all independent of $\ell$. In Appendix C we compute the subleading correction and show that it does not affect the asymptotic behavior.
@@
The presence of the branch cut along the imaginary axis is evident given the difference between approaching the axis from the left ($\theta\to\frac{\pi}{2}^{+}$) or the right ($\theta\to\frac{\pi}{2}^{-}$). The exponential suppression $R(\omega)\sim e^{-2\pi R_{s}\omega}$ as $\omega\to\infty$ reflects that, at fixed angular momentum, high–frequency waves are almost entirely absorbed by the black hole. This behavior was first obtained in ben-israel_black_2017; Basha:2018bvi via a WKB approximation. Crucially, we show in Appendix C that the asymptotic behavior (118) persists at subleading order in the $|R_{s}\omega|\gg 1$ expansion.
@@
On the other hand, our high-frequency approximation in (117) does not recover all aspects of the black hole solution; specifically, the solution (117) yields QNMs that lie exactly on the imaginary axis and are missing their non-zero real part nollert_quasinormal_1993; andersson_asymptotic_1993, as was already noted in Andersson:1996cm. In any case, we expect that retaining subleading terms in the high-frequency approximation can recover the missing real part, which we find evidence of in Appendix C.
@@ 4.4 Analyticity for generic compact objects
In the long-wavelength limit, any compact object can be effectively treated as a point particle, with its dynamics captured by a worldline EFT, where finite-size effects are incorporated through an action of the form:
@@
$\delta_{\ell=0}\,(\omega)$ $=R_{s}\,\omega\left(\frac{1}{2\epsilon}+\frac{1}{2}\log{\frac{4\omega^{2}}{\bar{\mu}_{\mathrm{IR}^{2}}}}-\frac{1}{2}+\gamma_{E}+\frac{C_{0,0}}{4\pi R_{s}}\right)+R_{s}^{2}\omega\left[\left(\frac{11\pi}{12}+\frac{C_{0,0}}{4R_{s}}\right)\sqrt{\omega^{2}}+\frac{C_{0,1}}{4\pi R_{s}^{2}}i\omega\right]$
@@
Establishing the region of analyticity of $R(\omega)$ for a generic compact object is not possible beyond the low-frequency approximation where the worldline EFT applies. The combined requirements of reflection symmetry under $\omega\to-\omega^{*}$, as expressed in (7), and consistency with the phase-shift (4.4) for $\omega>0$, uniquely determine the analytic structure at low frequencies. In particular, all logarithmic contributions are supported along the entire imaginary axis, in exact analogy with the black hole case.
@@ 4.5 Effect of infrared regulators on analytic structure
In this section, we show how the analyticity results of Section 4.1 extend to the S-matrix defined with free asymptotic plane waves, which is formally an infrared-divergent object. [^22] This is relevant since perturbative scattering amplitude computations, including in the worldline EFT of section 4.4, are usually performed with asymptotically free particle states, where IR divergences are unavoidable. In particular, we will consider here hard-cutoff regularization.
In hard-cutoff regularization, the potential is set to zero after $r>R_{\text{IR}}$, where $R_{\text{IR}}$ is the hard cutoff. Since the potential is zero at large $r$, the solution reduces to plane waves and one can define an S-matrix ${R}^{\text{hard}}(\omega)$ in the usual way Taylor:1972. Now, we can compare this hard-cutoff definition with the tortoise coordinate definition (28) where the potential falls-off fast enough. We find the relation
@@
for some finite $N$. This requirement excludes exponentially growing amplitudes and it is a key ingredient to ensure the validity of dispersion relations and causality constraints. [^24] In QFT, unitarity on the real axis, together with analyticity in the UHP and polynomial boundedness, implies that the partial wave $S_{\ell}(\omega)$ remains bounded along any direction in the UHP, $|S_{\ell}(s)|\leq 1$ for $|s|\to\infty$. Moreover, in the presence of a mass gap, one expects $|S_{\ell}(s)|\to 1$ asymptotically, since the effective interaction region becomes finite. This behavior is recovered by setting $R_{\text{IR}}=R_{s}$, where the S-matrix reduces to the tortoise-coordinate definition and approaches a constant in the UHP. Requiring polynomial growth on the regularized reflection coefficient $R^{\text{hard}}(\omega)$ imposes a sharp lower bound on the infrared cutoff:
@@ 5 Discussion and future directions
    Verification through explicit solutions. In Chapters 3 and 4, we verified the previously proven analyticity domains in exactly solvable cases. First, we considered scattering off a Pöschl–Teller potential, which can be seen as a toy model for the gravitational potential (see Figure 7). Then, in Chapter 4, we analyzed the Regge–Wheeler equation both at small frequencies, where the branch cuts associated with gravitational tails become manifest, and at high frequencies, where the asymptotic quasinormal-mode spectrum in the lower-half-plane can also be observed (see Figures 10 and 11 and Eqs. (107) and (117)).
@@
Physical origin of the branch cut in the upper half-plane. In QFT, singularities are typically associated with particle exchange, but this interpretation only applies partially here. We can associate some of the branch-cuts seen in the low-frequency approximation of the phase-shifts in Section 4.4 with Feynman diagrams. In particular, the single-graviton exchange and ladder diagrams (associated with the $1/r$ Newtonian potential) generate the $\log(\omega^{2})$ branch cut while triangle and higher “fan” diagrams, suppressed at larger distances, produce square-root-type cuts $\sqrt{\omega^{2}}$ Duff:1973zz; Caron-Huot:2025tlq:
@@
Interestingly, decreasing the cutoff below the Schwarzschild radius, $R_{\text{IR}}<R_{s}$, leads to exponential growth of the regulated amplitude for $\mathrm{Im}\,\omega\to+\infty$. This violate the standard assumption of polynomial boundedness in gapped S-matrix theory Correia:2020xtr. Requiring this assumption to hold leads to a lower bound on the IR cutoff,
@@
It would be interesting to see if such a bound holds more generally, beyond this classical wave setting, and in dimensional regularization.
@@
A closely related example is scattering of electromagnetic radiation on charged or polarizable materials, where no long-range interactions are present. In this case, the corresponding photon amplitudes are expected to be analytic in the full upper half-plane, making them suitable for a bootstrap analysis. Such an approach could in principle constrain electromagnetic properties of compact or point-like objects, such as their electric polarizabilities Jackson:1998nia. In Appendix D we illustrate analytically how such an analysis would proceed using tools from two-dimensional S-matrix bootstrap EliasMiro:2019kyf; EliasMiro:2021nul.
@@
Analyticity in Anti-de Sitter. The techniques used here to establish the analyticity of the S-matrix could, in principle, be extended to non–asymptotically flat settings, such as Anti–de Sitter (AdS) space. In that case, one would similarly need to define appropriate Jost solutions in terms of the free AdS modes. However, we expect that the upper half-plane would remain analytic for all connection coefficients, even in the presence of long-range interactions in the bulk. This is because the branch cut observed in flat space arises from Stokes phenomena, which in turn originate from the presence of an irregular singular point in the differential equation—something that typically does not occur in AdS Castro:2013lba; Denef:2009yy.
@@
It is conceivable that similar Stokes phenomena relating correlators and amplitudes also occur in QFT. While the retarded correlator should remain analytic in the upper half-plane of frequency space, the amplitude itself can develop non-analyticities associated with virtual particle exchange, analogous to the branch cut observed here for the elastic partial amplitude. Investigating this connection, even perturbatively, could be highly informative.

### v2

@@ 3.2 Inspection of the exact solution
As required by causality, the Green’s function is analytic in the UHP of $\omega$ for finite $x$ and $x^{\prime}$. However, as anticipated by our general proof of analyticity, $R(\omega)$ has poles on the positive imaginary axis where $A_{\mathrm{out}}(\omega)$ diverges, due to the term $\Gamma\left(\frac{i\omega}{\alpha}\right)$. Finally, we confirm that both $G_{R}(\omega,x,x^{\prime})$ and $R(\omega)$ are analytic within the strip $\alpha<\text{Im}\,\omega<\alpha$ due to the exponential decay of the potential (77) (as shown in Appendix A).
@@
We first notice that $\phi_{L}(\omega,x)$ in (78) is manifestly analytic for $\mathrm{Im}\,\omega>0$ because $_{2}F_{1}(a,b,c,\xi)$ only has singularities when $c$ is a non-positive integer. To see how the singularities of $R(\omega)$ appear, we rewrite the hypergeometric function using the identity
@@ 4.1 Analyticity domains of the S-matrix
Importantly, the spherical symmetry of the Schwarzschild background reduces the problem to one-dimensional scattering of waves in the radial direction $r$. However, since the gravitational potential behaves like $1/r$ at large distances, the asymptotic solutions are not given in terms of plane waves, and the S-matrix cannot be defined in terms of coefficients of outgoing/ingoing plane waves. Instead, the radial solutions reduce to Coulomb-type wavefunctions, which behave like
@@
the potential decays faster than $1/|x|$, as we show later, and the asymptotic states reduce to plane waves, allowing us to define an S-matrix as in (4). [^12] An analogous IR-safe definition exists for electromagnetic (Coulomb) scattering in relativistic quantum mechanics Lippstreu:2025jit.
On the opposite end, the tortoise coordinate maps the event horizon $r=R_{s}$ to $x=$-$\infty$, where waves can escape, leaving the system. Wave scattering on the Schwarzschild background therefore amounts to scattering on the real line across a potential barrier with open boundary conditions at both ends. This map allows us to rigorously apply the results of Section 2.
@@
Figure 9: The Schwarzschild potential barrier $V_{\text{RW}}(r(x))$ in units $R_{s}=1$ with $\ell=1,s=0$ for Left: $x\in\mathbb{R}$ and Right: $x\in\mathbb{C}$. The potential has an infinite number of poles on the imaginary axis, located at $x=i(1+2n)\pi R_{s},\;\;n\in\mathbb{Z}$. Each pole is a branch point, with the branch cuts chosen to run to $i\infty$ ($i\infty$) for branch points in the UHP (LHP). The potential decays exponentially as $x\rightarrow$-$\infty$ and decays as $x^{-2}$ as $x\to\infty$. The solid line on the right plot is the real axis, and the dashed line is slated by an angle $\theta$. We can take $\theta\in(-{\pi\over 2},{\pi\over 2})$ without encountering singularities.
@@
The Regge–Wheeler potential in the complex $x$-plane is shown in the right panel of Figure 9, confirming the analytic structure discussed above. [^14] Note that the region of the imaginary axis between $i\pi\,R_{s}$ and $i\pi\,R_{s}$ is free of branch cuts, so the potential remains analytic along this strip. The Zerilli potential exhibits the same analytic structure on the first sheet and is therefore not shown here.
@@
Moreover, since the potentials are strictly positive, all poles lie in the lower half of the complex frequency plane, $\mathrm{Im}\,\omega<0$ (see Section 2.6). These poles correspond to quasinormal modes and contribute to the retarded Green’s function, as well as to the transmission and reflection amplitudes.
@@ 4.2 Low-frequency expansion
where the real part of the phase-shift $\delta_{\ell}(\omega)$ captures elastic scattering and the imaginary part encodes absorptive effects. The phase shifts for $\ell=0,1$, obtained from the MST method Mano:1996vt; Mano:1996gn; Mano:1996mf, and adjusted for consistency with the reflection symmetry, Eq. (7), and the analytic domain derived above, take the form
@@ 4.3 Inspection of solvable cases
Newtonian potential. Following the derivation in liu_spectrum_1996; Andersson:1996cm, in the large-$r$ limit the Regge–Wheeler potential in (95) can be approximated as
@@
The solutions of Eq. (102) are the well-known Coulomb wavefunctions. Rather than constructing the retarded Green’s function from the Jost solutions $\phi_{L}(\omega,r(x))$ and $\phi_{R}(\omega,r(x))$, we instead employ $\widetilde{\phi}_{L}(\omega,r)$, defined as the solution that is regular and vanishes at the origin and $\widetilde{\phi}_{R}(\omega,r)\equiv\phi_{R}(\omega,r)$, which still represents an outgoing plane wave in the limit $x\to\infty$. Applying these boundary conditions, the solutions are given by
@@
and the reflection amplitude is given by, via (2), [^18] The amplitude (107) exhibits a series of poles on the imaginary axis at $\omega=-in/R_{s}$, with $n\in\mathbb{N}$. Such poles have appeared in various contexts (see, e.g., Kabat:1992tb; Bautista:2021wfy; Bautista:2023sdf), but they arise only within this approximation and are not present in the quasinormal-mode spectrum of the full S-matrix.
@@
The retarded Green’s function is analytic in the UHP of $\omega$, since the confluent hypergeometric functions are analytic for the relevant parameter ranges. This is shown explicitly in the left panel of Figure 10. Furthermore, $\widetilde{\phi}_{L}(\omega,r)$ is meromorphic everywhere (with poles in LHP), while $\widetilde{\phi}_{R}(\omega,r)$, and in turn $G_{R}(\omega,r,r^{\prime})$, has a branch cut along the negative imaginary axis of $\omega$. The discontinuity across this branch cut contributes to the late-time power-law decay tail of the black hole ringdown Leaver:1986gd; price_nonspherical_1972_2439; price_nonspherical_1972_2419; Casals:2012ng; Casals:2015nja. [^19] This can be seen by computing the inverse Fourier transform of the retarded Green’s function and deforming the integration contour into the LHP. Picking up the discontinuity of the branch cut results in the power-law behavior of $G_{R}(t\to\infty,r,r^{\prime})\sim t^{-n}$.
Figure 10: Solution to Newtonian potential in units $R_{s}=1$ for $\ell=1$. Dotted lines mark the real and imaginary axes. Left: $G_{R}(\omega,r=2,r^{\prime}=2.5)$. Right: $R(\omega)$. The Green’s function is analytic in the UHP, with QNMs (poles) appearing in the LHP, which lie on the imaginary axis. The reflection amplitude has QNMs in the LHP as well, but also has a branch cut along the entire imaginary axis arising from the Stokes phenomenon.
@@
The presence and logarithmic nature of the singularity are consistent with the low-frequency approximation of the previous section, where such logarithms are also present. Additionally, the location of the branch cut was anticipated from the analyticity argument of Section 2. We will see this same structure in the high-frequency approximation.
High-frequency approximation. We now approximate the Regge-Wheeler equation (92) in the high-frequency limit $|R_{s}\omega|\gg 1$. Again following the derivation in liu_spectrum_1996; Andersson:1996cm, the Regge-Wheeler equation can be written as
@@
whose solutions are known. The centrifugal term in the potential is subleading in this limit, so at this order the solutions are all independent of $\ell$. In Appendix C we also consider the subleading correction.
@@
The presence of the branch cut along the imaginary axis is evident given the difference between approaching the axis from the left ($\theta\to\frac{\pi}{2}^{+}$) or the right ($\theta\to\frac{\pi}{2}^{-}$). The exponential suppression $R(\omega)\sim e^{-2\pi R_{s}\omega}$ as $\omega\to\infty$ reflects that, at fixed angular momentum, high–frequency waves are almost entirely absorbed by the black hole. This behavior was first obtained in ben-israel_black_2017; Basha:2018bvi via a WKB approximation. Crucially, we also show here, in Appendix C, that the asymptotic behavior (118) persists at subleading order in the $|R_{s}\omega|\gg 1$ expansion.
@@
On the other hand, our high-frequency approximation in (117) does not quantitatively recover all aspects of the black hole S-matrix; specifically, our solution (117) yields QNMs that lie exactly on the imaginary axis. They miss a non-zero real part which has been established in nollert_quasinormal_1993; andersson_asymptotic_1993.232323The absence of real part of the QNMs was also observed in the high-frequency approximation of Andersson:1996cm. In any case, we expect that retaining subleading terms in the high-frequency approximation can recover the missing real part, which we find evidence of in Appendix C.
@@ 4.4 Analyticity for generic compact objects
In the long-wavelength limit, any compact object can be effectively treated as a point particle, with its dynamics captured by a worldline EFT, where finite-size effects are incorporated through an action of the form: [^24] This wordline EFT is for a scalar wave, see Combaluzier--Szteinsznaider:2025eoc for the case of gravitational perturbations.
@@
$\delta_{\ell=0}\,(\omega)$ $=R_{s}\,\omega\left(\frac{1}{2\epsilon}+\frac{1}{2}\log\bigg({\frac{4\omega^{2}}{\bar{\mu}_{\mathrm{IR}^{2}}}}\bigg)-\frac{1}{2}+\gamma_{E}+\frac{C_{0,0}}{4\pi R_{s}}\right)+R_{s}^{2}\omega\left[\left(\frac{11\pi}{12}+\frac{C_{0,0}}{4R_{s}}\right)\sqrt{\omega^{2}}+\frac{C_{0,1}}{4\pi R_{s}^{2}}i\omega\right]$
@@
Establishing the region of analyticity of $R(\omega)$ for a generic compact object is not possible beyond the low-frequency approximation where the worldline EFT applies. The combined requirements of reflection symmetry under $\omega\to-\omega^{*}$, as expressed in (7), and consistency with the phase-shift expression (4.4) for $\omega>0$, uniquely determine the analytic structure at low frequencies. In particular, all logarithmic contributions are supported along the entire imaginary axis, in exact analogy with the black hole case.
@@ 4.5 Effect of infrared regulators on analytic structure
In this section, we show how the analyticity results of Section 4.1 extend to the S-matrix defined with free asymptotic plane waves, which is formally an infrared-divergent object and must be regularized. [^25] This exercise is relevant since perturbative scattering amplitude computations, including in the worldline EFT of section 4.4, are usually performed with asymptotically free particle states, where IR divergences are unavoidable. We will consider here hard-cutoff regularization.
In hard-cutoff regularization, the potential is set to zero after some $r>R_{\text{IR}}$, where $R_{\text{IR}}$ is the hard cutoff. Since the potential is zero at large $r$, the solution reduces to plane waves and one can define an S-matrix ${R}^{\text{hard}}(\omega)$ in the usual way Taylor:1972. Now, we can compare this hard-cutoff definition with the tortoise coordinate definition (28) where the potential falls-off fast enough. We find the relation
@@
for some finite $N$. This requirement excludes exponentially growing amplitudes and it is a key ingredient to ensure the validity of dispersion relations and causality constraints. [^27] In QFT, unitarity on the real axis, together with analyticity in the UHP and polynomial boundedness, implies that the partial wave $S_{\ell}(\omega)$ remains bounded along any direction in the UHP, $|S_{\ell}(s)|\leq 1$ for $|s|\to\infty$. Moreover, in the presence of a mass gap, one expects $|S_{\ell}(s)|\to 1$ asymptotically, since the effective interaction region becomes finite. This behavior is recovered in our case by setting $R_{\text{IR}}=R_{s}$, where the S-matrix reduces to the tortoise-coordinate definition and approaches a constant in the UHP, as dictated by (118). Requiring polynomial growth on the regularized reflection coefficient $R^{\text{hard}}(\omega)$ imposes a sharp lower bound on the infrared cutoff:
@@ 5 Discussion and future directions
    Verification through explicit solutions. In Chapters 3 and 4, we verified the previously proven analyticity domains in exactly solvable cases. First, we considered scattering off a Pöschl–Teller potential, which can be seen as a toy model for the gravitational potential (see Figure 7). Then, in Chapter 4, we analyzed the Regge–Wheeler equation both at small frequencies, where the branch cuts associated with gravitational tails become manifest (see Eq. (99)), and at high frequencies, where the asymptotic quasinormal-mode spectrum in the lower-half-plane can also be observed (see Figure 11 and Eq. (117)). We also analyzed the exactly solvable Newtonian case where the analytic structure is manifest (see Figure 10 and Eq. (107)).
@@
Physical origin of the branch cut in the upper half-plane. In QFT, singularities are typically associated with particle exchange, but this interpretation only applies partially here. We can associate some of the branch-cuts seen in the low-frequency approximation of the phase-shifts in Section 4.4 with Feynman diagrams. In particular, the single-graviton exchange and ladder diagrams (associated with the $1/r$ Newtonian potential) generate the $\log(\omega^{2})$ branch cut while triangle and higher “fan” diagrams, suppressed at larger distances Duff:1973zz; Ivanov:2024sds; Caron-Huot:2025tlq, produce square-root-type cuts $\sqrt{\omega^{2}}$:
@@
Interestingly, decreasing the cutoff below the Schwarzschild radius, $R_{\text{IR}}<R_{s}$, leads to exponential growth of the regulated amplitude for $\mathrm{Im}\,\omega\to+\infty$. This violates the standard assumption of polynomial boundedness in gapped S-matrix theory Correia:2020xtr. Requiring this assumption to hold leads to a lower bound on the IR cutoff,
@@
It would be interesting to see if analogous bounds hold more generally in quantum field theories with massless particles.
@@
A closely related example is scattering of electromagnetic radiation on charged or polarizable materials, where no long-range interactions are present. In this case, the corresponding photon elastic amplitudes are expected to be analytic in the full upper half-plane, making them suitable for a bootstrap analysis. Such an approach could in principle constrain electromagnetic properties of compact or point-like objects, such as their electric polarizabilities Jackson:1998nia. In Appendix D we illustrate analytically how such an analysis would proceed using tools from the two-dimensional S-matrix bootstrap Paulos:2016but; EliasMiro:2019kyf; EliasMiro:2021nul.
@@
Analyticity in Anti-de Sitter and non-asymptotically flat backgrounds. The techniques used here to establish the analyticity of the S-matrix could, in principle, be extended to non–asymptotically flat settings, such as Anti–de Sitter (AdS) space. In that case, one would similarly need to define appropriate Jost solutions in terms of the free AdS modes. However, we expect that the upper half-plane would remain analytic for all connection coefficients, even in the presence of long-range interactions in the bulk. This is because the branch cut observed in flat space arises from Stokes phenomena, which in turn originate from the presence of an irregular singular point in the differential equation—something that typically does not occur in AdS Castro:2013lba; Denef:2009yy.
@@
It is conceivable that similar Stokes phenomena between correlators and amplitudes also occur in QFT. While the retarded correlator should remain analytic in the upper half-plane of frequency space, the amplitude itself can develop non-analyticities associated with virtual particle exchange, analogous to the branch cut observed here for the elastic partial amplitude. Investigating this connection, even perturbatively, could be highly informative.